OpenStudy (anonymous):

Determine if 1/2-4 is a solution to 4x-3y= -10 . Please show all of your work

OpenStudy (anonymous):

do you mean (1/2, -4)? If so... sub (1/2, -4) into the equation left side: 4(1/2) - 3(-4) = 2 +12 = 14 left side does not equal right side, so it is not a solution

OpenStudy (anonymous):

yes